What is the oxidation number of Mn in KMnO4 and of S in Na2S2O3?

Grade:12th pass

1 Answers

ankit singh
3 years ago

Let the Ox.no. of Mn in KMnO4 be x.
We know that, OX.no. of K=+1
Ox.no. of O=2
So, Ox. no. K+ Ox. no. Mn + 4(Ox.no.O)=0
or +1+x+4(2)=0
or +1+x8=0
or x=+81=+7
Hence, Ox.no. of Mn in KMnO4 is +7.
Similarly, for S in Na2S2O3,
2(Ox.no.Na)+2(Ox.no.S)+3(OX.no.O)=0
2×(+1)+2x+3(2)=0
x=+2
Hence, Ox. no. of S in Na2S2O3=+2.